A man deposits an amount of Rx into an investment account paying 8,76% p.a. compounded quarterly. Thirty years later he withdraws R50000 from the accumulated amount in the account and immediately invests the remainder in a retirement annuity which earns 9% p.a. The retirement annuity will provide him with a regular monthly income of R25000 for the next fifteen years starting one month after the investment is made. The original amount Rx (to the nearest rand) is equal to
1
Expert's answer
2020-05-21T17:49:16-0400

Solution: Find Rx 


Find the present value of the Annuity. 


Pv A= Annuity{1-(1+r)-n}/r


Pv A= 25000{1-(1+0.09÷12)-15*12}/0.09÷12


Pv A=25000*{1-(1+0.0075)-180}/0.0075


Pv A= 25000(98.5934)


Pv A=R2464835


Thus future value of the investment= R 2464835+R 50000


Future value of the investment=R 2514835


To find the original amount find the present value of rhe investment.


Pv=FV*{1÷(1+r)n}


Pv=FV/(1+r)n


Pv=2514835÷[(1+0.0876÷4)30*4]


Pv=2514835÷(1.0219)120



Pv=2514835÷13.4588


Pv=186853.66


Rx to the nearest Rand= 186854


=R 186,854